# Graph Reconstruction with a Connected Components Oracle

### v1

@@ 5 Relationship with the Maximal Independent Set Oracle
6 Relationship with the Separation Oracle
In this section, we compare the CC oracle against the separation oracle used in machine learning. We also discuss implications of known results for the separation oracle.
@@ 6 Relationship with the Separation Oracle
It is trivial to solve the GR problem with $\mathcal{O}(n^{2})$ Sep queries. We just query $\operatorname{Sep}(v,w,V\setminus\{v,w\})$ for each pair of vertices $v,w\in V,v\neq w$.
@@ 7 Conclusions
We proposed a novel connected components oracle and showed that it takes only a logarithmic number of queries to solve graph reconstruction in graphs of bounded number of edges, maximum degree, or treewidth. These results were complemented by unconditional lower bounds and a comparison against the MIS oracle.
@@
Another open question is to study whether a weaker oracle would yield similar upper bounds. For example, a weaker oracle could answer the following query: Given a set $Q$, does $G[Q]$ have one connected component?

### v2

@@ 5.2 Relationship with the Maximal Independent Set Oracle
5.3 Relationship with the Separation Oracle
Finally, we compare the CC oracle against the separation oracle used in machine learning. We also discuss implications of known results for the separation oracle.
@@ 5.3 Relationship with the Separation Oracle
It is trivial to solve the GR problem with $\mathcal{O}(n^{2})$ Sep queries. We just perform queries $\operatorname{Sep}(v,w,V\setminus\{v,w\})$ for each pair of vertices $v,w\in V,v\neq w$.
@@ 6 Conclusions
We proposed a novel connected components oracle and showed that it requires only a (poly)logarithmic number of queries to solve graph reconstruction in graphs of bounded number of edges, maximum degree, treewidth, or degeneracy. These results were complemented by unconditional lower bounds and comparisons against the #\#$C$ oracle, the MIS oracle, and the $\operatorname{Sep}$ oracle.
@@
Another direction is to study how much weaker the CC oracle could be made such that we would still obtain similar upper bounds. For example, an oracle that might be weaker than the CC oracle but stronger than the #\#$C$ oracle would be one that could answer the following query: Given a set $Q$, output one vertex from each connected component of $G[Q]$.